STS-130: EVA-3 marks Cupola’s grand opening – Comm issues evaluated

by Chris Bergin

International Space Station (ISS) crews have begun to enjoy a unique view of the planet below them, as STS-130’s spacewalkers removed the MLI blanket insulation blankets from the Cupola’s windows. The final spacewalk of the mission saw Bob Behnken and Nick Patrick perform a six hour EVA, as engineers on the ground review the cause of communication issues on the first two spacewalks.

STS-130 Latest:

Most of the mission’s major milestones had been completed by the time EVA-3 started, with Node 3 now being outfitted since its installation, which was soon followed by the relocation of the Cupola to the Earth-facing port, and the Pressurized Mating Adaptor (PMA-3) move to the port side of Tranquility.

“STS-130 mission is proceeding nominally. Flight Day 8/9 Activities Completed: Cupola vestibule outfitting was started. Pressurized Mating Adapter-3 (PMA-3) un-berthed from Harmony zenith port and installed on forward port of Tranquility (Node 3),” summed up the mission status via the NASA Test Director (NTD).

“Flight: Day 9/10 Activities Planned: EVA-3 procedure review and campout by Behnken and Patrick. EVA-3 begins Tuesday at 2109 hrs EST. Node 3 to PMA 3 reconfiguration of fluid lines and cables.

The Mission Evaluation Room (MER) continues to enjoy a smooth mission, thanks to the performance of Endeavour – with no issues recorded since docking. As is normal for a “quiet” mission, engineers have been spending time evaluating a few of the minor issues recorded earlier in the flight.

“MER Items: MER-06: TCS (Trajectory Control Sensor) erratic data during rendezvous,” was one previous “funny” (minor issue) reviewed – which holds no mission impact for undocking later this week.

“Problem Description: During approach, TCS did not operate as expected. TCS did not transition from the Pulsed to Continuous Wave (CW) laser mode as expected at ~1000 feet. Instead TCS remained with the Pulsed laser and had occasional instances of switching to the CW laser,” listed MMT documentation via L2.

“During the RPM (Range ~600 feet), the CW laser data became ratty and usable. The raw Rdot values were fluctuating between 0.7 and 5 feet per second, with the filter smoothing the data enough to be usable. The Pulsed laser was overridden by the crew on the V-bar (Range ~250 feet).

“The CW laser performance on the V-bar was no better than the earlier Pulse laser performance, showing jumps in position, rapid changes in Rdot and frequent nav-initializations. No further troubleshooting was performed.

STS-130 Specific Articles:

With no TCS data available from the RPM onwards towards docking, the crew used the HHL (Hand Held Laser) and range ruler as the prime sensors and/or backup TRAD sensor for docking. The HHL will be used for the post-undocking flyaround.

“For undocking and fly-around, TCS expects that the CW laser will continue to generate unusable data with jumps in range/range rate that was observed during rendezvous,” added the presentation, before adding the most probably cause for the TCS issue.

“Analysis of the TCS data for the rendezvous and docking portion of the mission indicated that the nonconformance was being generated from the intermediate tone board. This board is used to generate the range value for the CW returned signal. A review of the STS-127 data for the previous flight of SN 1007 did not show this irregularity. The spikes in the range value coincided with an intermittent failure of a bit in the affected board.”


Behnken and Patrick concluded the final EVA of the mission – their third in a row – following on the heels of two very successful spacewalks. Most of the work was conducted around the new hardware that flew uphill with them on Endeavour.

“After the standard campout protocol is completed the crew presses into a 6.5 hr EVA. The crew perform the STS/ISS safing for the EVA (Inhibit Pad),” noted EVA-3 procedures (L2).

“The first task is to install the PMA-3 heater and data cables. Followed by the opening of the Loop B QDs (Quick Disconnect) on Node 3 and removal of the LTA (Launch To Activation) cables.”

No strangers to the MLI blanket insulation – following their work to install numerous blankets to protect Node 3’s ammonia lines on EVA-2 – the duo were tasked with the removal of the MLI from the windows on Cupola. Storing the bulky material into large bags was challenging, but successful.

“The next big task is to remove all of the MLI on the external face of the Cupola. A bag will be setup to hold all of it, to be brought inside and down to the ground. The launch locks will be release once they are done with the MLI,” pre-empted the procedures.

Once the launch locks were removed, the ISS crew cycled the window shields/shutters one at a time, providing them with the first view of the Earth from their new observation deck. All of the windows weren’t open at the same time, with Tuesday’s task simply used to check the shutters opened without a problem.

A few hours later, all of the windows were opened together, providing a stunning view for the the crewmembers – see live thread for FD10 for all the EVA’s events, with screenshots and video.

With Patrick already completing the task of installing most of the handrails on Node 3 as a previous EVA get-ahead, only a few additional tasks were added to the end of the spacewalk. Most of this work will aid STS-131’s planned spacewalks.

These get-ahead tasks included the relocating an APFR (Articulated Portable Foot Restraint), work on a VSC Video Cable – stretching back to the Russian segment – and work on a gap spanner.

EVA Comm issue evaluations:

Controllers on the ground closely monitoring the quality of the communications between the spacewalkers during the spacewalk, the ISS and the ground – following some minor issues during both EVA-1 and EVA-2.

“MER-07: EV1 CCA SN 1213 volume low during EVA 1. MER-08: EV1 CCA SN 1172 No C&W tones during EVA 2 prep,” lists the two issue reports at the MER – with most of the evaluation work concentrating on EVA-1’s problem.

“Problem Description: EVA1: During STS-130 EVA-1, EV1 (Behnken) earphone/receive audio in both earphones began to degrade approximately 4 hours into the EVA. EV1 increased the volume from 4 to 6. He continued the EVA with low audio levels; not total loss of audio.

“Prior to entering the Airlock at the end of the EVA, EV1 switched from PRI (primary) to ALT (alternative) with no change in audio level. Post EVA the ear cups were inspected for visible moisture with none found. Seals were inspected; seals looked intact. A comm check was performed in both hard-line and PRI using the Prime CCA and the B/U (Back UP) CCA. The B/U had nominal audio with the Prime still exhibiting the same reduced audio level.

The problem is related to the CCA – or Comm Cap – worn by spacewalkers during their EVAs – with the likely cause expected to relate to a build-up of moisture, even though it wasn’t obvious via post-EVA inspections. Communication problems were also noted on two EVAs during the previous shuttle mission, STS-129.

“History: STS-129 EV1, EMU 3018 on EVA 1 – Low audio (max vol) during last 2 hrs EVA; post EVA, dried CCA ear cups. STS-129 EV1, EMU 3018 on EVA 2 – EVA Prep: good comm check and tones. EVA: reported only 1 alert tone, faint comm (max vol) – Post-EVA Troubleshooting: low audio (tones, comm) compared to different CCA,” outlined a dedicated MMT presentation (L2).

“STS-130 EV1, EMU 3018 on EVA 1 – Low audio with (max vol) during last hour of EVA. Post EVA tested B/U CCA and reported good audio (comm only). STS-130 EV1, EMU 3018 on EVA 2 – EVA Prep: B/U CCA no tones; switched to EV2 b/u CCA (good check out). No audio issues during EVA.

“Concern: Post STS-129, CCA team investigating audio issues. EMU 3018 is common in each of these EVA audio/comm issues. EVA Team Response/Action. EVA Team with FCE and EV met to evaluate the audio/comm system. Developed fault tree (high-level) and event timeline. No quick answer, many system variables limited data – 4 electrical modules, 6 separate harnesses, multiple CCAs, multiple anomaly signatures.”

With EMU 3018 set for return onboard Endeavour, engineers will be able to examine all the associated hardware in order to confirm root cause and potential get-wells for future EVAs. No communication issues were noted during EVA-3.

L2 members : Documentation – from which the above article has quoted snippets – is available in full in the related L2 sections, now over 4500 gbs in size

Related Articles